At the end of the year, the records of NCIS Corporation provided the following selected and

incomplete data:

  • Common shares: $10 par value at the beginning of the year

  • Shares authorised: 200,000

  • Shares issued: all shares were issued at $17 per shares. Total cash collected:

    $2,125,000

  • There was a 2:1 stock split during the year.

  • Treasury stock: 3,000 shares (repurchased at $20 per share).

  • Treasury stock acquired after the stock split.

  • Net income $240,340.

  • Dividends declared and paid $123,220.

  • Retained earnings balance at the beginning of the year: $555,000.

    Required:

  1. Determine the number of authorised, issued and outstanding shares at the end of the year.

  2. What is the balance of the share premium account?

  3. What was the dividend paid per share?

  4. In what section of the balance sheet should the treasury stock be reported and for what

    amount?

  5. What is the par value per share at the end of the year?

  6. Provide, if necessary, the journal entry to record the stock split. If no journal is required

    explain why.
Answer #1

authorised shares = 200000

issued shares = 2125000 / 17 =125000

outstanding shares =( (125000/2)*3 ) - 3000 =184500

balance in share premium account = (2125000/17)*7 - (3000*13.3) = 835100

dividend per share = 123220 / 184500 = 0.668 per share

the cost of treasury stock is reported at the end of the Stockholders' Equity section of the balance sheet. Treasury stock will be a deduction from the amounts in Stockholders' Equity. amount = 3000*20 = 60000

par value = 2shares*$10 = 20 / 3shares = $6.7

journal entry for recording the stock split

account & title debit credit
share capital ($10 par) 1250000
share capital (6.7 par value) 1250000